OpenStudy (crashonce):

When the expression mx^2 + 6x +n is divisible by x-3, find the value of 18m+2n-3

OpenStudy (crashonce):

@jim_thompson5910 @ganeshie8

OpenStudy (mathmath333):

\(mx^2 + 6x +n\) is divisible by \((x-3)\) hence \(\large\tt \begin{align} \color{black}{f(3)=m(3)^2+6(3)+n=0}\end{align}\)

OpenStudy (crashonce):

@mathmath333 continure

OpenStudy (mathmath333):

did u simplify that

OpenStudy (mathmath333):

\(\large\tt \begin{align} \color{black}{9m+n+18=0}\end{align}\)

OpenStudy (mathmath333):

try to multiply it by \(2\)

OpenStudy (crashonce):

18m + 2n+36=0

OpenStudy (crashonce):

so the final answer t the question is -39

OpenStudy (mathmath333):

\(\large\tt \begin{align} \color{black}{9m+n+18=0\\ 9m +n=-18\\18m+2n=-36\\18m+2n-3=-36-3=-39}\end{align}\)

OpenStudy (crashonce):

thanks